# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

### v1

@@ 3 Data
The analysis is based on the latest wave of the OECD PISA (Programme for International Student Assessment) for the year 2022 which measures secondary students’ learning achievement (at 15 years-old), i.e., their ability to use their reading, mathematics and science knowledge and skills to meet real-life challenges.1 The paper is focused on the LAC region, which includes a sample of 10 countries, namely: Argentina, Brazil, Chile, Colombia, Dominican Rep., Mexico, Panama, Peru, Paraguay and Uruguay.2 Since the aim of the paper is about identifying the determinants of learning poverty (i.e., those students located at the bottom of the learning distribution in three subjects: math, reading and science) in the region, I employ two comparisons: no competency achieved (level 0 which includes level1b and level1c) versus level 1 (level1a), and level 1 versus level 2 (where those below level 2 are referred to as ‘low performers’; OECD, 2023b, p. 89). These two working samples contain, respectively, between 16,236 and 9,484 observations.
It should be mentioned that the PISA 2022 data gathering was mediated by the pandemic. Originally, it was planned to take place in 2021 but the disruption caused by COVID-19 global pandemic forced the assessment to be postponed by a year, to 2022, whose content was updated with further questions linking the “resilience of education systems” from the point of view of fairness and inclusion within the pandemic’s remote learning (OECD, 2023a).
3.1 Data coding—rationale
My focus on the three bottom levels of achievement is aligned to the observed adverse long-term learning trajectory of the region where, for example, the rate of low-performers in math among the poorest students either increased or remained unchanged across all countries since 2012 (Arias Ortiz et al., 2024a, b). As said above, I use two binary educational outcomes (or dependent variables) among the low performance group which are cardinally defined towards those students achieving the least: level 0 (=1) versus level 1 (=0), and level 1 (=1) versus level 2 (=0). As shown in Table 1, these two comparisons//samples capture most of the LAC students. For instance, levels 0 to 2 (first three rows of Table 1) account for 89.5% (math), 77.5% (reading) and 80.9% (science) of the total students.
@@ 3.1 Data coding—rationale
Regarding data processing, rather than analysing low achievement drivers for specific subjects as in other PISA empirical ML studies (see: Gómez-Talal et al., 2024; Huang et al., 2024, and the systematic review: Wang et al., 2023), I follow a subject-aggregated definition of dependent variables –for instance, $y_{\text{level 0,level1}}$ = 1 if a student falls into level 0 for math, reading and science and equals to 0 if a student falls into level 1 for the three subjects. This leads to rates for the first educational outcome (levels 0 versus 1) of 65.3% and for the second educational outcome (level 1 versus level 2) of 59.3% (Panel B of Table 1). I constructed two datasets for each educational outcome to be analysed separately: a dataset with student//family variables and a dataset with school variables. Besides, after selecting leading covariates with lower missing rates, I eliminated records with missing values (e.g., $\approx$ 3% in the student//family and school datasets).3 Full details on covariates construction and definitions are shown in Appendix C (Tables C1 and C2).
@@ 3.2.1 Students and family covariates
The empirical relevance of students and family factors as drivers of the probability of falling into the sub-groups of bottom performers (level 0 versus level 1) and low performers (level 1 versus level 2) can be seen by the summary statistics of Table 2. On the one hand, in comparison to students whose achievement fall into level 1, level 0 (no achieving) students (Table 2, columns 1 to 3), are likely to do less homework, have a lower number of digital devices (= -1.72) and fewer books (= -0.35), being twice as likely to be from an immigrant family and three time as likely to speak another language at home (Panel A, columns 1 to 3). Repetition for the level 0 group of students —either at primary or lower secondary— is much larger (33% against 9% and 22% against 9%) as well as the chances of missing or skipping school as the intensity of paid work (1.66 versus 1.45). Further, students from the level 0 group compared to students from level 1 group are poorer (69% comes from families in the bottom half of the family SES index distribution in comparison to 59% in the case of level 1 students) and come from larger families (Panel B, column 3, Table 2).
@@
On the other hand, these gaps on covariates are similar and hold when comparing level 1 versus level 2 students (Table 2, columns 4 to 6), although with some differences in magnitude with the level 0-level 1 comparison. For instance, the gap on possession of books at home and chances of paid work are larger, but repetition gaps are smaller. Notably, the gaps on parental education become relevant (and statistically significant) as well as the importance of family educational support.
@@ 3.2.2 School covariates
As shown by Table 3, school context —be its composition, inputs and infrastructure, teachers profile or policies implemented— matters when related to the chances of students falling into either the lower achieving group level 0 (among bottom performers, columns 1 to 3) or level 1 (among low performers, columns 4 to 6). The influence of school context is slightly more prominent when comparing the bottom two levels of performance. Here, I mention some key differences. When contrasting to students who achieve in level 1, level 0’s students are more likely to come from rural schools (by 13%) and public schools (by 12%), and from schools where there is less competition, which are located in smaller towns, which received less funding from fees and which are smaller in terms of enrolment (Table 3, column 3). Likewise, students at the bottom of the learning distribution attend school where there are less PC for teaching connected to the web and with less involvement of parents and with fewer methods for students’ testing.
@@
In addition, the school prevalence of disadvantage students now becomes an important factor between level 1 and level 2 groups, being 9% higher for the latter group (column 6, Table 3). The remaining covariate raw differences are, to a large extent, similar to the previous students’ comparison. There are some distinct pattern in covariates differences, however. For example, no statistical differences are found for STRs, the proportion of certified teachers, infrastructure, and in some school policies (e.g., transfer, improvement policies, and ability grouping). On the whole, summary statistics from Tables 2 and 3 underline who contextual disadvantages stemming from student, family and school factors are vital for the likelihood of students falling into the lowest performing groups in the LAC region.
@@ 4.1 Binary classification models
Given the dichotomous nature of dependent variables, I employ five binary classification ML models (for details of the five ML models below, see: Hastie et al., 2009, 2015 and for its implementation: Pedregosa et al., 2011; Raschka and Mirjalili, 2019), for the two dichotomous dependent variables (level 0-level 1 and level 1-level 2). Each model was trained using the stratified $K$-fold (5 folds) method and train-test split of 80%-20% of the data so that to ensure the data is balanced within each fold within this undersampling strategy. Each model was optimised using different hyperparameters setup and the application of grid search to find the top performing model version among each type of model. Below, I briefly describe each of the five ML models employed.
@@ 4.1.1 Logistic regression
Logistic regression (or just logit) is one the principal models for binary outcomes, which predicts the probability that the outcome equals to one conditional on a vector of covariates and the error term following a logistic distribution:
$Pr\big(y_{i}=1|\mathbf{x}\big)=\text{exp}\big({\beta_{0}+\mathbf{x}_{i}\boldsymbol{\beta}}^{\prime}\big)/\big(1+\text{exp}\big({\beta_{0}+\mathbf{x}_{i}\boldsymbol{\beta}}^{\prime}\big)$ (1)
Logit is simpler and easier to interpret. Here, the hyperparameters used for the two versions are lack of regularisation (no penalty), and with penalty L2 (Ridge) option.
@@ 4.1.2 Logistic Lasso with CV penalty
I also use a different version of logistic regression; namely, a Logistic Lasso (with cross-validation, CV) (logitlasso) which imposes a penalty (L1) on the size of coefficients to control overfitting and the size of it is chosen by CV. The lasso penalises the absolute size of coefficients via the penalty parameter $\lambda$, yielding sparse solutions where some coefficient estimates are set to exactly zero. I used four values for $C$ (the inverse of regularization parameter values) for cross-validation ($C$ = 10, 50, 100 and 200). The penalised objective function for lasso is:
$Q_{\mathcal{L}}=\sum_{i=1}^{N}\widetilde{w}_{i}\times\mathcal{L}+\lambda\sum_{j=1}^{P}\kappa_{j}|\beta_{j}|$ (2)
where $\mathcal{L}$ is the logit likelihood function.
@@ 4.1.3 Random forest
Random Forest (RF) improves modelling by ensembling a combination of multiple decision trees and then aggregating their predictions. RF trains several decision tree models (learners) separately and then let all the learners decide the output of the test samples –the final classification– by a voting mechanism. That is, the predicted class is: $\widehat{\mathbf{y}}=\text{mode}\big\{y_{1},y_{2},...,y_{n}\big\}$, and $\mathbf{y}_{i}$ is the prediction of the $i$th tree. Also, the training samples of each learner are randomly chosen and when building the decision tree a portion of the features (variable) is randomly selected from the features to build the decision tree. The hyperparameters assessed are the number of estimators (or trees), with 100, 500 and 1,000 trees chosen and two maximum depths of each tree given by 5 and 10 leaves.
@@ 4.1.4 Gradient boosting
Another ML classifier method I use is Gradient Boosting (GB) under six formats (or combination of hyperparameters): the number of trees ranging from 100, 500 and 1,000 and two learning rates (1% and 10%). GB is an ensemble of predicted decision trees models run in a sequential fashion correcting prediction of previous models, and the update rule is:
$f_{m}=f_{m-1}(\mathbf{x})+\sum_{j=1}^{J_{m}}\gamma_{jm}1(\mathbf{x}\in R_{jm})$ (3)
where $f_{m}$ is the model $m$ interaction and $J$ are terminal nodes, $\gamma_{jm}$ the fitted values and $R_{jm}$ the residuals.
@@ 4.1.5 Neural networks
Neural nets (NN) or Feed-forward neural networks consist of hidden layers that link the predictors (or input layers) to the outcome. Each hidden layer is composed of multiple units (nodes) that pass signals to the next layer using an activation function. Central tuning parameters are the choice of the activation function and the number and size of hidden layers. I consider six specifications for the NN: one neural net with one hidden layers of 200 nodes each and another neural net with three hidden layer set of 200, 100 and 50 nodes (200, 100, 50), using three activation rules.
@@ 4.2 Assessment of binary classification models
I assess the performance of each of different ML model versions (made by the configuration of hyperparameters) relying on an array of evaluation metrics (Naidu et al., 2023). Specifically, I use six metrics when comparing the models: Accuracy (ACC), Recall (RC), Precision (PR), F1 Score (F1), Specificity (SP) and the Area Under the Curve (AUC). These metrics provides overall measures of model performance. Noteworthy is the AUC indicator because, by calculating the area of under the Receiver Operating Characteristic (ROC) curve, it can efficiently balance the true positive rate and the false positive rate. I include a description of the six metrics in Appendix 0.A.
@@ 4.3 Stacking approach for classification
When faced with a classification task as in the current paper (and after choosing the best model based on evaluation metrics) the next decision is to assess which one is best suited. But this not obvious, and combining multiple learners (ML algorithms) into a final ensemble model, is likely to lead to superior performance compared with each individual learner. Stacked generalisation (or stacking), a concept introduced by Wolpert (1992), allows a form of model averaging using a technique that utilises the predictions of various models to train a meta-model. Theoretically, stacking perform as well as individual learners (Van der Laan et al., 2007). Here, I follow the stacking approach among the best five binary classification ML models using the work of Ahrens et al. (2023), implementing stacking in Stata.
@@
The stacking process for binary classification exercises consists of two steps. First, cross-validated predicted probabilities for each of the five base learners and each observation $\big(\widehat{y}^{-k(i)}_{(j),i}\big)$ are obtained for all folds ($K$ data partitions of similar size) expect from the fold where current observation is located (i.e., k-k(ii)). Second, the final learner (stacked model) is fit by using the observed $y_{i}$ outcome and cross-validated predicted probabilities as predictors $\big(\widehat{y}^{-k(i)}_{(1),i},\widehat{y}^{-k(i)}_{(2),i},...,\widehat{y}^{-k(i)}_{(J),i}\big)$, with the condition that weights are zero or positive ($w_{i}\geq$ 0) and the sum of weights are equal to one $\sum_{j=1}^{J=5}w_{j}=1$, where $j$ denotes different base learners. That is, for the standard least squares:
$min_{w_{1},...,w_{J}}\sum_{i=1}^{N}\Big(y_{i}-\underbrace{\sum_{j=1}^{J}w_{j}\widehat{y}^{-k(i)}_{(j),i}}_{\widehat{y}_{i}^{*}}\Big)^{2}$ (4)
@@
where $\widehat{y}_{i}^{*}$ are the stacking predicted probabilities, a weighted sum of predicted probabilities from rerunning learner $j$ on the full sample $I$.
4.4 SHAP values—interpretable ML estimates
In the body of educational research using ML (for a review, see: Ersozlu et al., 2024; Hilbert et al., 2021), there has been a latest growth in the application of Shapley values to identify main features behind educational performance in educational surveys (for instance, Gómez-Talal et al., 2024; Huang et al., 2024). Here I define the key concepts related to Shapley values (a more detailed and technical description of Shapley values can be found in Appendix 0.B). Importantly, I rely on SHAP values to present leading estimates in Section 5.

### v2

@@ 3 Data
The analysis is based on the latest wave of PISA for the year 2022 which measures secondary students’ learning achievement (at 15 years-old), i.e., their ability to use their reading, mathematics and science knowledge and skills to meet real-life challenges.2 The paper is focused on the LAC region, which includes a sample of 10 countries, namely: Argentina, Brazil, Chile, Colombia, Dominican Rep., Mexico, Panama, Peru, Paraguay and Uruguay.3 Since the paper’s aim is to identify the determinants of learning poverty (i.e., those students located at the bottom of the learning distribution in three subjects: math, reading and science) in the region, we employ two comparisons: no competency achieved (level 0 which includes level1b and level1c) versus level 1 (level1a), and level 1 versus level 2 (where those below level 2 are referred to as ‘low performers’; OECD, 2023b, p. 89). These two working samples contain, respectively, between 16,236 and 9,484 observations.
It should be mentioned that PISA 2022 data gathering was mediated by the pandemic. Originally, it was planned to take place in 2021 but the disruption caused by COVID-19 global pandemic forced the assessment to be postponed by a year (to 2022) and the content was updated with further questions linking the “resilience of education systems” from the point of view of fairness and inclusion within the remote learning shift brought about by the pandemic (OECD, 2023a). The resilience of education systems in PISA 2022 was evaluated against a new framework focusing on students’ capacity for continuous learning and adaptability. In addition to the test measuring students’ academic performance, there is a comprehensive questionnaire that collects relevant information about the students’ family and socioeconomic backgrounds, their attitudes toward learning and study habits and schools’ background features. Data submission in the LAC region happened in December 2023.
3.1 Data coding–rationale
The paper’s focus on the three bottom levels of achievement is aligned to the adverse long-term learning trajectory of the region where, for example, the rate of low-performers in math among the poorest students either increased or remained unchanged across all countries since 2012 (Arias Ortiz et al., 2024a, b). As earlier mentioned, we use two binary educational outcomes (or dependent variables) among the low performance groups: level 0 (=1) versus level 1 (=0), and level 1 (=1) versus level 2 (=0). As shown in Table 1, these two samples capture most of the LAC students. For instance, levels 0 to 2 (first three rows of Table 1) account for 89.5% (math), 77.5% (reading) and 80.9% (science) of total students.
@@ 3.1 Data coding–rationale
Regarding data processing, rather than analysing low achievement drivers for specific subjects as in other PISA empirical ML studies (see: Gómez-Talal et al., 2024; Huang et al., 2024, and the systematic review: Wang et al., 2023), the paper follows a subject-aggregated definition of dependent variables–for instance, $y_{\text{level 0,level1}}$ = 1 if a student falls into level 0 for math, reading and science, and equals to 0 if a student falls into level 1 for the three subjects. This leads to rates for the first educational outcome (level 0 versus level 1) of 65.3% and for the second educational outcome (level 1 versus level 2) of 59.3% (Panel B of Table 1). We construct two datasets for each educational outcome to be analysed separately: a dataset with student//family variables and a dataset with school variables. Besides, after selecting leading covariates with lower missing rates, we eliminate records with missing values (e.g., $\approx$ 3% in the student//family and school datasets).4 Full details on covariates’ construction and definitions are shown in the Appendix (Tables C1 and C2).
@@ 3.2.1 Students and family covariates
The empirical relevance of students and family covariates as determinants of probabilities of students to fall into the sub-groups of bottom performers (level 0 versus level 1) and low performers (level 1 versus level 2) can be seen by the summary statistics of Table 2. On the one hand, in comparison to students whose achievement fall into level 1, level 0 students (Table 2, columns 1 to 3), are likely to do less homework, have a lower number of digital devices (= -1.72) and fewer books (= -0.35), being twice as likely to be from an immigrant family and three time as likely to speak another language at home (Panel A, columns 1 to 3). Repetition for the level 0 group of students –either at primary or lower secondary–is much larger (33% against 9% and 22% against 9%, respectively) and likewise the chances of missing or skipping school, and their intensity to be engaged in paid work (1.66 versus 1.45). Further, students from the level 0 group compared to students from level 1 group are poorer (69% comes from families in the bottom half of the family SES index distribution in comparison to 59% in the case of level 1 students) and come from larger families (Panel B, column 3, Table 2). On the other hand, these gaps on covariates are similar and hold when comparing level 1 versus level 2 students (Table 2, columns 4 to 6), although with some differences in magnitude with the level 0-level 1 comparison. For instance, the gap on possession of books at home and chances of paid work are larger, but repetition gaps are smaller. Notably, gaps on parental education become relevant (and statistically significant) along with the importance of family educational support.
@@ 3.2.2 School covariates
As shown by Table 3, school’s context–be its composition, inputs and infrastructure, teachers profile or policies implemented–matters when it comes down to a student’s chance of falling into either the lower achieving group level 0 (among bottom performers, columns 1 to 3) or level 1 (among low performers, columns 4 to 6). The influence of school’s context is slightly more prominent when comparing the bottom two levels of performance. Here, we mention some key differences. Compared to students whose achievement falls into the level 1 performance bracket, level 0’s students are more likely to come from rural schools (by 13%) and public schools (by 12%), and also from schools where there is less competition with other schools, and from schools which are located in smaller towns and receive less funding from fees, and schools which are smaller in terms of enrolment (Table 3, column 3). What’s more, students at the bottom of the learning distribution attend schools where there are less PC per teacher connected to the web, and with less involvement of parents and with fewer methods for students’ testing.
@@
Besides, the school rate of disadvantage students now becomes an important factor between level 1 and level 2 groups, being 9% higher for the latter group (column 6, Table 3). The remaining covariates raw differences are, to a large extent, similar to the previous students’ comparison. There are some distinct pattern in covariates differences, however. For example, no statistical differences are found for the student-teacher ratio (STR), the proportion of certified teachers, infrastructure, and for some school policies (e.g., transfer, improvement policies, and ability grouping). On the whole, summary statistics from Tables 2 and 3 underline to what degree contextual disadvantages stemming from student, family and school levels are critical for the likelihood of students falling into the lowest performing groups.
@@ 4.1 Binary classification models
Given the dichotomous nature of dependent variables (level 0-level 1 and level 1-level 2), we employ five binary classification ML models (for details see: Hastie et al., 2009, 2015 and for their implementation: Pedregosa et al., 2011; Raschka and Mirjalili, 2019). Each model was trained using the stratified $K$-fold (5 folds) method and train-test split of 80%-20% of the data so that to ensure the data is balanced within each fold. Each model was optimised using different hyperparameters setup and the application of grid search to find the top performing model version among each type of model. Below, we briefly describe each ML model employed.
@@ 4.1.1 Logistic regression
Logistic regression (or just logit) is one the principal models for binary outcomes. It predicts the probability that the outcome equals to one conditional on a vector of covariates:
$\text{Pr}\big(y_{i}=1|\mathbf{x}\big)=\Lambda(\mathbf{x}^{\prime}\boldsymbol{\beta})\equiv\text{exp}\big({\beta_{0}+\mathbf{x}^{\prime}_{i}\boldsymbol{\beta}}\big)/\big(1+\text{exp}\big({\beta_{0}+\mathbf{x}^{\prime}_{i}\boldsymbol{\beta}}\big)$ (1)
where $\Lambda(.)$ is c.d.f. of the logistic distribution. Logit is simpler and easier to interpret. Here, the hyperparameters used for the two versions are lack of regularisation (no penalty), and with penalty L2 (Ridge) option. The likelihood function for logit is:
$\mathcal{L}=\sum_{i=1}^{N}\big[y_{i}(\beta_{0}+\mathbf{x}^{\prime}_{i}\boldsymbol{\beta})-\text{log}\big(1+e^{(\beta_{0}+\mathbf{x}^{\prime}_{i}\boldsymbol{\beta})}\big)\big]$ (2)
@@ 4.1.2 Logistic Lasso with CV penalty
We also rely on a different version of logistic regression; namely, a Logistic Lasso with cross-validation (CV) (or logitlasso), which imposes a penalty (L1) on the size of coefficients to control overfitting and the size of it is chosen by CV. The lasso penalises the absolute size of coefficients via the penalty parameter $\lambda$, yielding sparse solutions where some coefficient estimates are set to exactly zero. We use four values of $C$ (the inverse of regularization strength) for CV ($C$ = 10, 50, 100 and 200). The penalised objective function for logitlasso combines the standard logistic regression loss function (negative log-likelihood) with a L1 penalty term for the coefficients (the second term):
$Q_{\mathcal{L}}=-\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}+\lambda\sum_{j=1}^{p}|\beta_{j}|$ (3)
where $\mathcal{L}$ is the logit likelihood function, $\lambda$ is a non-negative tuning parameter that controls the strength of the regularisation, and $p$ is the number of predictor variables (excluding the intercept).
@@ 4.1.3 Random forest
Random Forest (RF) improves modelling by ensembling a combination of multiple decision trees and then aggregating their predictions. RF trains several decision tree models (learners) separately and then let all the learners decide the output of the test samples (i.e., the final classification) by a voting mechanism. The final binary classification prediction is a majority vote across all the individual decision trees. Formally:
$\widehat{\mathbf{y}}=\text{mode}\big[f_{b}(\mathbf{x}^{\prime})\big]_{b=1}^{B}$ (4)
where $\widehat{\mathbf{y}}$ is the final predicted binary class (bottom or low performer = 1, and 0 otherwise), $B$ denotes the total number of trees, $f_{b}(\mathbf{x}^{\prime})$ is the prediction for the input $\mathbf{x}^{\prime}$ for the $b$-th tree. Moreover, the training samples of each learner are randomly chosen and when building the decision tree a portion of the features (variables) is randomly selected from the features to build the decision tree. The hyperparameters configuration we assess are the number of estimators (or trees): 100, 500 and 1,000 trees, and each tree’s depth: two maximum depths given by 5 and 10 leaves.
@@ 4.1.4 Gradient boosting
Another ML classifier method we use is Gradient Boosting (GB) which builds decision trees iteratively to minimize prediction errors. GB constructs weak learners (trees) sequentially, and each tree predicts the residual of earlier predictions. That is, GB is an ensemble of predicted decision trees models run in a sequential fashion correcting prediction of previous models, and the update rule is:
$f_{m}=f_{m-1}(\mathbf{x})+\sum_{j=1}^{J_{m}}\gamma_{jm}\mathbf{1}(\mathbf{x}\in R_{jm})$ (5)
where $f_{m}$ is the model $m$ interaction and $J$ are terminal nodes, $\gamma_{jm}$ the fitted values and $R_{jm}$ the residuals. We assess the GB under six formats (or combination of hyperparameters): the number of trees ranging from 100, 500 and 1,000 and two learning rates (1% and 10%).
@@ 4.1.5 Neural networks
The final ML model we consider is Neural nets (NN) or Feed-forward neural networks. It consists of hidden layers that link the predictors (or input layers) to the outcome. Each hidden layer is composed of multiple units (nodes) that pass signals to the next layer using an activation function. Central tuning parameters are the choice of the activation function and the number and size of hidden layers. We consider six specifications: one NN with one hidden layer of 200 nodes each, and another neural net with three hidden layers with 200, 100 and 50 nodes, and both using three activation rules (i.e., logistic, relu, tanh).
@@ 4.2 Assessment of binary classification models
We assess the performance of each ML model version by relying on an array of evaluation metrics (Naidu et al., 2023). Specifically, six metrics are employed when comparing models: Accuracy (ACC), Recall (RC), Precision (PR), F1 Score (F1), Specificity (SP) and the Area Under the Curve (AUC). These metrics provide overall measures of a model’s performance. Noteworthy is the AUC indicator because, by calculating the area of under the Receiver Operating Characteristic (ROC) curve, it can efficiently balance the true positive rate and the false positive rate. We include a description of the six metrics in Appendix 0.A.
@@ 4.3 Stacking approach for classification
When faced with a classification task as in the current paper–and after choosing the best model configuration for each of the five ML models based on evaluation metrics–the next decision is to assess which model is best suited. But this decision is not obvious and, combining multiple learners into a final ensemble model, is quite likely to lead to a superior performance compared with each individual learner. Ensemble models represent a paradigm in which multiple algorithms are strategically combined to improve prediction accuracy and robustness beyond what can be achieved by any single model (Qiao and Hu, 2020). Our choice of the five types of ML classification models to build up the ensemble model is due to their diversity, so that errors or better fittings across models are complemented. For instance, logit and logitlasso are linear models whereas RF and GB are tree-based ensemble methods, and NN can account for model highly complex, non-linear relationships. This choice is somehow validated by systematic reviews on the topic of ML for the identification of students at risk where ensemble models combining RF, GB and SVM are more commonly employed (Shi et al., 2025).
@@
Stacked generalisation (or stacking), a concept introduced by Wolpert (1992), allows a form of model averaging using a technique that utilises the predictions of various models to train a meta-model. Theoretically, stacking performs as well as individual learners (Van der Laan et al., 2007). Here, we follow the stacking approach among the best five binary classification ML models following the work of Ahrens et al. (2023).5 The stacking process for binary classification exercises consists of two steps.6 First, cross-validated predicted probabilities for each of the five base learners and each observation $\big(\widehat{y}^{-k(i)}_{(j),i}\big)$ are obtained for all folds ($K$ data partitions of similar size), expect from the fold where the current observation is located (i.e., k-k(ii)). Second, the final learner (stacked model) is fit by using the observed $y_{i}$ outcome and cross-validated predicted probabilities as predictors $\big(\widehat{y}^{-k(i)}_{(1),i},\widehat{y}^{-k(i)}_{(2),i},...,\widehat{y}^{-k(i)}_{(J),i}\big)$ of the full sample. The final learner is obtained by minimising the least squares with the additional constraint that coefficients sum to one:
$\begin{split}\underset{w_{1},...,w_{J}}{\text{min}}&=\sum_{i=1}^{N}\Big[(y_{i}-\underbrace{\sum_{j=1}^{J}w_{j}\widehat{y}^{-k(i)}_{(j),i}}_{\widehat{y}_{i}^{*}}\Big]^{2}\\ \text{s.t. }&w_{j}\geq 0\text{ and }\sum_{j=1}^{J}w_{j}=1\end{split}$ (6)
@@
where here $J$ = 5 (the number of models//learners), stacking predicted probabilities are defined as $\widehat{y}_{i}^{*}=\sum_{j}\widehat{w}_{j}\widehat{y}_{(j),i}$ and $\widehat{w}_{j}$ is the estimated stacking weight for learner $j$, and $\widehat{y}_{(j),i}$ are the predicted probabilities when refitting learner $j$ on the full sample.
4.4 SHAP values–interpretable ML estimates